8. Determine the first three nonzero terms in the Taylor approximation for y + cos y = e, y(0) = 0, 7(0) = 1.
Answer #1

g(0)=0 Yozd y = ex-cosy y co= 1- Cosco) = 1-1=0. y - ett siny. y yCO)= 1 +) = 1. yive ext(cosy.y+ys Griny) = y yo = 1+